OpenStudy (anonymous):

graph the area bounded by f(x) = (-x^2 ) 4x 1 and g(x) = x 1

OpenStudy (anonymous):

we have to first find the limits of integration

OpenStudy (anonymous):

set both function equal to each other and solve

OpenStudy (anonymous):

What is the function? Is it -x^2+4x+1 or -4x+1 or -4x-1? Is it x-1 or x+1?

OpenStudy (anonymous):

i have no idea how to use this site

OpenStudy (anonymous):

it -x^2+4x+1

OpenStudy (anonymous):

is it x+1 or x-1?

OpenStudy (dumbcow):

malevolence i dont think they mean find the area under curve...graph the area, like shading the bounded area this is graphing quadratic and linear functions not integration i believe. nelson is this for calculus?

OpenStudy (anonymous):

yea Dumbcow calc 1 i have a lot more questions if you want to help me

OpenStudy (anonymous):

I can help you with anything calculus involved. I just need to know the functions we're dealing with. Is it x+1 or x-1?

OpenStudy (anonymous):

the question is graph the area bounded by f(x) = (-x^2 ) 4x +1 and g(x) = x +1

OpenStudy (anonymous):

then calculate the area

OpenStudy (anonymous):

Okay. You need to set the two equations equal to find the bounds for you integral: -x^2+4x+1=x+1 -x^2+3x=0 x^2-3x=0 x(x-3)=0 So your bounds are 0,3 Now you need to figure out which curve is on top. Plug in 2 and see. (-4+8+1)=5, 2+1=3 So you have the integral of the top curve minus bottom curve. \[\int\limits_0^3 (-x^2+4x+1)-(x+1)dx\] Combine like terms. Integrate using power rule. Can you finish it?

OpenStudy (dumbcow):

ok i was wrong...follow malevolence steps

OpenStudy (anonymous):

no i don't know anything

OpenStudy (anonymous):

Okay, well combining: \[\int\limits_0^3(-x^2+3x)dx\] Then you want to use the power rule that says: \[\int\limits x^n dx=\frac{x^{n+1}}{n+1}\] So: \[\frac{-x^3}{3}+\frac{3x^2}{2}|_0^3=\frac{-1}{3}(3^3-0^3)+\frac{3}{2}(3^2-0)=\frac{-27}{3}+\frac{27}{2}=\frac{9}{2}\]